# Large Language Model for Generative Recommendation

### v1

@@ 3.1. Sequence Generation
Methods movielens 25M Toys HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10 P5 0.0688 0.0464 0.1040 0.0577 $\bm{0.0239}$ $\bm{0.0145}$ $\bm{0.0411}$ $\bm{0.0201}$ GenRec $\bm{0.1034}$ $\bm{0.0716}$ $\bm{0.1311}$ $\bm{0.0837}$ 0.0190 0.0136 0.0251 0.0157
@@ 4.1. Dataset
We conduct extensive experiments on two real-world Amazon (Ni et al., 2019) datasets and MovieLens (Harper and Konstan, 2015) dataset to evaluate the performance of our proposed GenRec approach on recommendation tasks. The Amazon datasets, which record user purchase histories across a diverse range of products, were sourced from the Amazon.com platform. MovieLens datasets comprise a large number of movie ratings and associated metadata, contributed by users of the MovieLens website over various periods. The descriptive statistics of these datasets are depicted in Table 1 (see reference 2). For each user interaction sequence, the most recent item is used as the test data, the second-most recent is used as validation data, and the remaining is used for training.
Dataset Movielens 25M Toys

### v2

# GenRec: Large Language Model for Generative Recommendation

@@ 3.1. Sequence Generation
Methods MovieLens 25M Amazon Toys HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10 P5 0.0688 0.0464 0.1040 0.0577 $\bm{0.0239}$ $\bm{0.0145}$ $\bm{0.0411}$ $\bm{0.0201}$ GenRec $\bm{0.1034}$ $\bm{0.0716}$ $\bm{0.1311}$ $\bm{0.0837}$ 0.0190 0.0136 0.0251 0.0157
@@ 4.1. Dataset
We conduct extensive experiments on two real-world datasets from Amazon (Ni et al., 2019) and MovieLens (Harper and Konstan, 2015), respectively, to evaluate the performance of our proposed GenRec approach on recommendation tasks. The Amazon datasets, which record user purchase histories across a diverse range of products, were sourced from the Amazon.com platform. MovieLens datasets comprise a large number of movie ratings and associated metadata, contributed by users of the MovieLens website over various periods. The descriptive statistics of these datasets are depicted in Table 1 (see reference 2). For each user interaction sequence, the most recent item is used as the test data, the second-most recent is used as validation data, and the remaining is used for training.
Dataset MovieLens 25M Amazon Toys